# TELEClass: Taxonomy Enrichment and LLM-Enhanced Hierarchical Text Classification with Minimal Supervision

### v2

@@ 3.4 Text Classifier Training with Path-Based Data Augmentation
4 Experiments
4.1 Experiment Setup
Datasets. We use two public datasets in different domains for evaluation. (1) Amazon-531 McAuley and Leskovec (2013) consists of Amazon product reviews and a three-layer label taxonomy of 531 product types, with 29,487 unlabeled training samples and 19,685 testing samples. (2) DBPedia-298 Lehmann et al. (2015) consists of Wikipedia articles with a three-layer label taxonomy of 298 categories, with 196,665 unlabeled training samples and 49,167 testing samples.
Compared Methods. We compare TELEClass with the following methods: Hier-0Shot-TC Yin et al. (2019), zero-shot ChatGPT (i.e., GPT-3.5-turbo), Hier-doc2vec Le and Mikolov (2014), WeSHClass Meng et al. (2019), TaxoClass Shen et al. (2021) and its variant TaxoClass-NoST that does not apply self-training. We also include a Fully-Supervised baseline that uses the entire labeled training data to train the text-matching network. See more details of these methods in Appex. B.1. The implementation details of TELEClass are in Appx. B.3.
Evaluation Metrics. Following previous studies Shen et al. (2021), we utilize the following evaluation metrics: Example-F1, Precision at k (P@k) for $k=1,3$, and Mean Reciprocal Rank (MRR). See Appendix B.2 for detailed definitions.
4.2 Experimental Results
Table 1 shows the evaluation results of all the compared methods. We make the following observations. (1) Overall, TELEClass achieves significantly better performance than other strong zero-shot and weakly-supervised baselines. (2) By comparing with other weakly-supervised methods, we find that TELEClass significantly outperforms TaxoClass-ST, the strongest baseline that, like TELEClass, does not use self-training. This further demonstrates the effectiveness of TELEClass. (3) Although LLMs (e.g., ChatGPT) show superior power in many tasks, naïvely prompting it in the hierarchical text classification task yields significantly inferior performance compared to strong weakly-supervised text classifiers. We conduct a more detailed comparison with LLM prompting for hierarchical text classification in Section 4.4.
Table 2: Performance of TELEClass and its ablations on Amazon-531 and DBPedia-298.
@@ 4.2 Experimental Results
Table 3: Performance comparison of TELEClass and zero-shot LLM prompting.
@@
F1 P@1 P@3 F1 P@1 P@3
ChatGPT 0.516 0.681 0.475 0.482 0.533 0.455
ChatGPT-L 0.662 0.857 0.644 0.665 0.830 0.649
GPT-4 0.699 0.822 0.689 0.605 0.652 0.592
TELEClass 0.648 0.851 0.642 0.863 0.935 0.863
4.3 Ablation Studies
@@
4.4 Comparison with Zero-Shot LLM Prompting
In this section, we further compare TELEClass with zero-shot LLM prompting. Because it is not straightforward to get ranked predictions by LLMs, we only report Example-F1 and P@k as performance evaluations. We report the zero-shot performance of ChatGPT and GPT-4 by including all the classes in the prompt and asking the LLM to provide the three most appropriate classes for a given document. Experiments on GPT-4 are conducted on 1000 randomly sampled test documents given the prohibitively high cost. We also include a level-by-level version, ChatGPT-L, which selects one most appropriate class at each level and continues to its children.
Table 3 shows the experiment results. We find that TELEClass consistently outperforms zero shot performance of LLMs by large margins on DBPedia-298. On Amazon-531, while it continues to show substantial improvements over ChatGPT, it slightly underperforms ChatGPT-L and GPT-4. As for the cost, once trained, TELEClass does not require an additional cost on inference and also has a substantially shorter inference time. Prompting LLMs takes a longer time and can be prohibitively expensive, and the cost will scale up with an increasing size of test data. The estimated API cost for ChatGPT on the entire DBPedia test set is \$80 and for GPT-4 it is around \$2,500. Also, we find that the level-by-level version ChatGPT-L consistently outperforms the naïve version, demonstrating the necessity of taxonomy structure.
Table 4: Intermediate results on two documents, including selected core classes by different methods, true labels, and the corresponding taxonomy enrichment results. The optimal core class in the true labels is marked with ©.
@@ 4.4 Comparison with Zero-Shot LLM Prompting
DBPedia The Lindenhurst Memorial Library (LML) is located in Lindenhurst, New York, and is one of the fifty six libraries that are part of the Suffolk Cooperative Library System … TaxoClass: village TELEClass initial: building TELEClass refined: library library©, agent, educational institution Class: library Enrichment: national library, central library, collection, volumes…
@@
Amazon Since mom (89 yrs young) isn’t steady on her feet, we have place these grab bars around the room. It gives her the stability and security she needs. TaxoClass: personal care, health personal care, safety TELEClass initial: daily living aids, medical supplies equipment, safety, TELEClass refined: bathroom aids safety health personal care, medical supplies equipment, bathroom aids safety© Class: bathroom aids safety Enrichment: seat, toilet, shower, safety, handles…
4.5 Case studies
@@
To better understand the TELEClass framework, we show some intermediate results of two documents in Table 4, including the core classes selected by (1) the original TaxoClass Shen et al. (2021) method, (2) TELEClass’s initial core classes selected by LLM, and (3) refined core classes of TELEClass. Besides, we also include the true labels of the documents and the taxonomy enrichment results of the corresponding core class in the table. Overall, we can see that TELEClass’s refined core class is the most accurate. For example, for the first Wikipedia article about a library, TaxoClass selects “village” as the core class, while TELEClass’s initial core class finds a closer one “building” thanks to the power of LLMs. Then, with the enriched classes-indicative features as guidance, TELEClass’s refined core class correctly identifies the optimal core class, which is “library”. In the other example, TELEClass also pinpoints the most accurate core class “bathroom aids safety” while other methods can only find more general or partially relevant classes.
5 Related Work
5.1 Weakly-Supervised Text Classification
@@
Weakly-supervised text classification trains a classifier with limited guidance, aiming to reduce human efforts while maintaining high proficiency. Various sources of weak supervision have been explored, including distant supervision Gabrilovich and Markovitch (2007); Chang et al. (2008); Song and Roth (2014) like knowledge bases, keywords Agichtein and Gravano (2000); Tao et al. (2018); Meng et al. (2018, 2020); Wang et al. (2021a); Zhang et al. (2021a), and heuristic rules Ratner et al. (2016); Badene et al. (2019); Shu et al. (2020). Later, extremely weakly-supervised methods are proposed to solely rely on class names to generate pseudo labels and train classifiers. LOTClass Meng et al. (2020) utilizes MLM-based PLM as a knowledge base for extracting class-indicative keywords. X-Class Wang et al. (2021a) extracts keywords for creating static class representations through clustering. PIEClass Zhang et al. (2023c) employs PLMs’ zero-shot prompting to obtain pseudo labels with noise-robust iterative ensemble training. MEGClass Kargupta et al. (2023) acquires contextualized sentence representations to capture topical information at the document level. WOTClass Wang et al. (2023) ranks class-indicative keywords from generated classes and extracts classes with overlapping keywords.
5.2 Hierarchical Text Classification
@@
A hierarchy provides a systematic top-to-down structure with inherent semantic relations that can assist in text classification. Typical hierarchical text classification can be categorized into two groups: local approaches and global approaches. Local approaches train multiple classifiers for each node or local structures Banerjee et al. (2019); Liu et al. (2005); Wehrmann et al. (2018). Global approaches, learn hierarchy structure into a single classifier through recursive regularization Gopal and Yang (2013), a graph neural network (GNN)-based encoder Peng et al. (2018); Huang et al. (2019); Zhou et al. (2020); Shen et al. (2021), or a joint document label embedding space Chen et al. (2021).
@@
Weak supervision is also studied for hierarchical text classification. WeSHClass Meng et al. (2019) uses a few keywords or example documents per class and pretrains classifiers with pseudo documents followed by self-training. TaxoClass Shen et al. (2021) follows the same setting as ours which uses the sole class name of each class as the only supervision. It identifies core classes for each document using a textual entailment model, which is then used to train a multi-label classifier. Additionally, MATCH Zhang et al. (2021c) and HiMeCat Zhang et al. (2021b) study how to integrate associated metadata into the label hierarchy for document categorization with weak supervision.
5.3 LLMs as Generators and Annotators
@@
Large language models (LLMs) have demonstrated impressive performance in many downstream tasks and are explored to help low-resource settings by synthesizing data as generators or annotators Ding et al. (2023). For data generation, few-shot examples Wang et al. (2021b) or class-conditioned prompts Meng et al. (2022) are explored for LLM generation and the generated data can be used as pseudo training data to further fine-tune a small model as the final classifier Ye et al. (2022). Recently, Yu et al. (2023) proposed an attribute-aware topical text classification method that incorporates ChatGPT to generate topic-dependent attributes and topic-independent attributes to reduce topic ambiguity and increase topic diversity for generation. For data annotation, previous works utilize LLMs for unsupervised annotation Feng et al. (2021), Chain-of-Thought annotation with explanation generation He et al. (2024), and active annotation Zhang et al. (2023a).
6 Conclusion and Future Work
@@
In this paper, we propose a new method, TELEClass, for the minimally-supervised hierarchical text classification task with two major contributions. First, we enrich the input label taxonomy with LLM-generated and corpus-based class-indicative terms for each class, which can serve as additional features to understand the classes and facilitate classification. Second, we explore the utilization of LLMs in the hierarchical text classification in two directions: data annotation and data creation. On two public datasets, TELEClass can outperform existing baseline methods substantially, and we further demonstrate its effectiveness through ablation studies. We also conduct a comparative analysis of performance and cost for zero-shot LLM prompting for the hierarchical text classification task. For future works, first, we plan to generalize TELEClass’s idea into other low-resource text mining tasks with hierarchical label spaces, such as fine-grained entity typing. We also plan to explore how to extend TELEClass into more complicated label spaces, such as an extremely large space with millions of labels.
Limitations
In this paper, we propose TELEClass, a novel hierarchical text classification method that uses the label taxonomy as the only supervision. Despite its effectiveness shown in the experiments, there is still room to improve. First, TELEClass utilizes the general knowledge of pretrained LLM and shows strong performance on two general domain datasets. However, it remains a challenge to train effective text classifiers in private domains, for which LLMs lack relevant knowledge. Second, in this paper, we mainly focus on acquiring high-quality pseudo labeled data while only utilizing the simplest classifier model and objective. It is worth studying how the proposed method can be further improved with more advanced network structure and noise-robust training objectives. Lastly, TELEClass assumes the availability of a large unlabeled corpus for pseudo labeling, whose performance is not fully studied if limited data is available.
Ethics Statement
In this work, we develop a new hierarchical text classification method and experiment on two publicly available datasets for research purposes. Overall, we do not foresee any significant ethical issues from our work. All the datasets are used as intended. We will open-source our code upon publication to assist future studies. Although showing strong language understanding capability, LLMs can sometimes generate misinformation Chen and Shu (2023) or biased contents Fang et al. (2024). In our work, instead of studying the generation ability of LLMs, we only use LLMs to generate pseudo documents for text classifier training. In practice, our method can be combined with any misinformation detection and debiasing methods Mattern et al. (2022); Chen and Shu (2021).
Acknowledgments

### v3

@@ 3.4. Text Classifier Training with Path-Based Data Augmentation
4. Experiments
4.1. Experiment Setup
4.1.1. Datasets
We use two public datasets in different domains for evaluation. Table 1 shows the data statistics.
  • Amazon-531 (McAuley and Leskovec, 2013) consists of Amazon product reviews and a three-layer label taxonomy of product types.
  • DBPedia-298 (Lehmann et al., 2015) consists of Wikipedia articles with a three-layer label taxonomy of its categories.
4.1.2. Compared Methods
  • Hier-0Shot-TC (Yin et al., 2019) is a zero-shot approach, which utilizes a pretrained textual entailment model to iterative find the most similar class at each level for a document.
  • GPT-3.5-turbo is a zero-shot approach that queries GPT-3.5-turbo by directly providing all classes in the prompt.
  • Hier-doc2vec (Le and Mikolov, 2014) is a weakly-supervised approach, which first trains document and class representations in a same embedding space, and then iteratively selects the most similar class at each level.
  • WeSHClass (Meng et al., 2019) is a weakly-supervised approach using a set of keywords for each class. It first generates pseudo documents to pretrain text classifiers and then performs self-training.
  • TaxoClass (Shen et al., 2021) is a weakly-supervised approach that only uses the class name of each class. It first uses a textual entailment model with a top-down search and corpus-level comparison to select core classes, which are then used as pseudo training data. We include both its full model, TaxoClass, and its variation TaxoClass-NoST that does not apply self-training on the trained classifier, which is the same as TELEClass.
  • TELEClass is our newly proposed weakly-supervised approach that only uses the class name for each class.
  • Fully-Supervised is a fully-supervised baseline that uses the entire labeled training data to train the text matching network used in TELEClass.
4.1.3. Evaluation Metrics
Following previous studies (Shen et al., 2021), we utilize the following evaluation metrics: Example-F1, Precision at k (P@k) for $k=1,3$, and Mean Reciprocal Rank (MRR). See Appendix C for detailed definitions.
4.1.4. Implementation Details
We use Sentence Transformer (Reimers and Gurevych, 2019) all-mpnet-base-v2 as the text encoder for the similarity measure in Section 3.1 and Section 3.3. We query GPT-3.5-turbo-0125 for LLM-based taxonomy enrichment, core class annotation, and path-based generation. For corpus-based taxonomy enrichment, we get the term and class name embeddings using a pre-trained BERT-base-uncased (Devlin et al., 2019), we select top $k=20$ enriched terms for each class (c.f. Section 3.2). We generate $q=5$ documents with path-based generation for each class. The document encoder in the final classifier is initialized with BERT-base-uncased for a fair comparison with the baselines. We train the classifier using AdamW optimizer with a learning rate 5e-5, and the batch size is 64. The experiments are run on one NVIDIA RTX A6000 GPU.
Table 3. Performance of TELEClass and its ablations on Amazon-531 and DBPedia-298 datasets. The best score is boldfaced.
@@ 4.1.4. Implementation Details
4.2. Experimental Results
Table 2 shows the evaluation results of all the compared methods. We make the following observations. (1) Overall, TELEClass achieves significantly better performance than other strong zero-shot and weakly-supervised baselines, which demonstrates the effectiveness of TELEClass on the hierarchical text classification task without any human supervision. (2) By comparing with other weakly-supervised methods, we find that TELEClass significantly outperforms TaxoClass-NoST, the strongest baseline that, like TELEClass, does not use self-training. Given that TELEClass uses an even simpler classifier model than TaxoClass-NoST, its superior performance shows the substantially better pseudo training data obtained by combining unlabeled corpus and LLMs. (3) Although LLMs (e.g., ChatGPT) show power in many tasks, naïvely prompting it in the hierarchical text classification task yields significantly inferior performance compared to strong weakly-supervised text classifiers. This proves the necessity of incorporating corpus-based knowledge to improve label taxonomy understanding for the hierarchical setting. We conduct a more detailed comparison with LLM prompting for hierarchical text classification in Section 4.4.
We also study the temporal complexity of TELEClass. We observe that, on Amazon-531, both TELEClass and the strongest baseline TaxoClass take around 5 to 5.5 hours. The reason why TELEClass does not increase the overall temporal complexity is that TaxoClass needs to run the textual entailment model on each pair of document and candidate class. On the other hand, the taxonomy enrichment step of TELEClass makes it possible to simplify this process with embedding similarity calculation which saves a lot of time, while the saved time is budgeted for LLM prompting.
Table 4. Performance comparison of TELEClass and zero-shot LLM prompting. We only report Example-F1 and P@k, because it is not straightforward to get ranking of classes predicted by LLMs for MRR calculation. We also report estimated costs in US dollars and running time in minutes for each method on the entire test set. “” indicates that we report the performance based on an estimation from a 1,000-document subset of test data.
@@ 4.2. Experimental Results
Example-F1 P@1 P@3 Est. Cost Est. Time Example-F1 P@1 P@3 Est. Cost Est. Time
GPT-3.5-turbo 0.5164 0.6807 0.4752 \$60 240 mins 0.4816 0.5328 0.4547 \$80 400 mins
GPT-3.5-turbo (level) 0.6621 0.8574 0.6444 \$20 800 mins 0.6649 0.8301 0.6488 \$60 1,000 mins
GPT-4 0.6994 0.8220 0.6890 \$800 400 mins 0.6054 0.6520 0.5920 \$2,500 1,000 mins
TELEClass 0.6483 0.8505 0.6421 <\$1 3 mins 0.8633 0.9351 0.8633 <\$1 7 mins
4.3. Ablation Studies
@@
4.4. Comparison with Zero-Shot LLM Prompting
In this section, we further compare TELEClass with zero-shot LLM prompting. Because it is not straightforward to get ranked predictions by LLMs, we only report Example-F1 and P@k as performance evaluation. Additionally, we report the estimated cost and time for each method on the entire test set. The inference time is reported in minutes, and please be aware that this is just a rough estimation as the actual running time is also dependent on the server condition.
We include the following settings to compare with TELEClass:
  • GPT-3.5-turbo: We include all the classes in the prompt and ask GPT-3.5-turbo model to provide 3 most appropriate classes for a given document.
  • GPT-3.5-turbo (level): We perform level-by-level prompting using GPT-3.5-turbo. Starting from the root node, we ask the model to return one most appropriate class for a given document, and we iteratively prompt the model with the children of the selected node at each level. This method can only generate a path in the taxonomy, but in the actual multi-label hierarchical classification setting, the true labels may not sit in the same path.
  • GPT-4: We include all the classes in the prompt and ask GPT-4 to provide 3 most appropriate classes for a given document. Given the limited budget, we only test it on randomly sampled 1,000 documents and estimate the cost on the entire test set.
Table 4 shows the experiment results. We find that TELEClass consistently outperforms all compared methods on DBPedia, while on Amazon, TELEClass underperforms GPT-3.5-turbo (level) and GPT-4 but still being comparable. As for the cost, once trained, TELEClass does not require additional cost on inference and also has substantially shorter inference time. Prompting LLMs takes longer time and can be prohibitively expensive (e.g., using GPT-4), and the cost will scale up with increasing size of test data. Also, we find that GPT-3.5-turbo (level) consistently outperforms the naïve version, demonstrating the necessity of taxonomy structure. It saves the cost because of the much shorter prompts, but takes longer time due to more queries made per document.
Table 5. Intermediate results on two documents, including selected core classes by different methods, true labels, and the corresponding taxonomy enrichment results. The optimal core class in the true labels is marked with ©.
@@ 4.4. Comparison with Zero-Shot LLM Prompting
DBPedia The Lindenhurst Memorial Library (LML) is located in Lindenhurst, New York, and is one of the fifty six libraries that are part of the Suffolk Cooperative Library System … TaxoClass: village TELEClass initial: building TELEClass refined: library library©, agent, educational institution Class: library Top Enrichment: national library, central library, collection, volumes…
@@
Amazon Since mom (89 yrs young) isn’t steady on her feet, we have placed these grab bars around the room. It gives her the stability and security she needs. TaxoClass: personal care, health personal care, safety TELEClass initial: daily living aids, medical supplies equipment, safety, TELEClass refined: bathroom aids safety health personal care, medical supplies equipment, bathroom aids safety© Class: bathroom aids safety Top Enrichment: seat, toilet, shower, safety, handles…
4.5. Case studies
@@
To better understand the TELEClass framework, we show some intermediate results of two documents in Table 5, including the core classes selected by (1) the original TaxoClass (Shen et al., 2021) method, (2) TELEClass’s initial core classes selected by LLM, and (3) refined core classes of TELEClass. Besides, we also include the true labels of the documents and the taxonomy enrichment results of the corresponding core class in the table. Overall, we can see that TELEClass’s refined core class is the most accurate. For example, for the first Wikipedia article about a library, TaxoClass selects “village” as the core class, while TELEClass’s initial core class finds a closer one “building” thanks to the power of LLMs. Then, with the enriched classes-indicative features as guidance, TELEClass’s refined core class correctly identifies the optimal core class, which is “library”. In the other example, TELEClass also pinpoints the most accurate core class “bathroom aids safety” while other methods can only find more general or partially relevant classes.
Besides, although TELEClass outperforms the zero-shot LLM prompting in most cases, there are cases showing the contrary and here is one example we found from Amazon-531. The product review is about “glycolic treatment pads” for which GPT correctly predicts its labels as “beauty” and “skin care”, while TELEClass predicts it as “health care”. We suspect that the word “treatment” in the review leads to the error because of the bias of term-based pseudo-labeling. It is a known issue of keyword-based methods and some solutions are proposed for the weakly-supervised flat text classification (Dong et al., 2023; Zhang et al., 2023a). We hope our study can motivate more research to solve this issue in the hierarchical setting.
5. Related Work
5.1. Weakly-Supervised Text Classification
@@
Weakly supervised text classification trains a classifier with limited guidance, aiming to reduce human efforts while maintaining high proficiency. Various sources of weak supervision have been explored, including distant supervision (Gabrilovich and Markovitch, 2007; Chang et al., 2008; Song and Roth, 2014) like knowledge bases, keywords (Agichtein and Gravano, 2000; Tao et al., 2018; Meng et al., 2018, 2020; Wang et al., 2021a; Zhang et al., 2021b), and heuristic rules (Ratner et al., 2016; Badene et al., 2019; Shu et al., 2020). Later, extremely weakly-supervised methods are proposed to solely rely on class names to generate pseudo labels and train classifiers. LOTClass (Meng et al., 2020) utilizes MLM-based PLM as a knowledge base for extracting class-indicative keywords. X-Class (Wang et al., 2021a) extracts keywords for creating static class representations through clustering. WDDC (Zeng et al., 2022) prompts a masked language model and train a classifier over the word distributions. NPPrompt (Zhao et al., 2023) uses the PLM embeddings to construct verbalizers and their weighted distribution over each class. PESCO (Wang et al., 2023a) performs zero-shot classification with semantic matching between class description and then fine-tune a classifier with iterative contrastive learning. PIEClass (Zhang et al., 2023a) employs PLMs’ zero-shot prompting to obtain pseudo labels with noise-robust iterative ensemble training. MEGClass (Kargupta et al., 2023) acquires contextualized sentence representations to capture topical information at the document level. WOTClass (Wang et al., 2023b) ranks class-indicative keywords from generated classes and extracts classes with overlapping keywords. RulePrompt (Li et al., 2024) mutually enhances automatically mined logical rules and pseudo labels for fine-tuning a classifier.
5.2. Hierarchical Text Classification
@@
A hierarchy provides a systematic top-to-down structure with inherent semantic relations that can assist in text classification. Typical hierarchical text classification can be categorized into two groups: local approaches and global approaches. Local approaches train multiple classifiers for each node or local structures (Banerjee et al., 2019; Liu et al., 2005; Wehrmann et al., 2018). Global approaches, learn hierarchy structure into a single classifier through recursive regularization (Gopal and Yang, 2013), a graph neural network (GNN)-based encoder (Peng et al., 2018; Huang et al., 2019; Zhou et al., 2020; Shen et al., 2021), or a joint document label embedding space (Chen et al., 2021). Recent studies also show that LLMs cannot comprehend the complex hierarchical structure (Bhambhoria et al., 2023; Fuhan et al., 2024; Yu et al., 2023a).
@@
Weak supervision is also studied for hierarchical text classification. WeSHClass (Meng et al., 2019) uses a few keywords or example documents per class and pretrains classifiers with pseudo documents followed by self-training. TaxoClass (Shen et al., 2021) follows the same setting as ours which uses the sole class name of each class as the only supervision. It identifies core classes for each document using a textual entailment model, which is then used to train a multi-label classifier. Additionally, MATCH (Zhang et al., 2021c) and HiMeCat (Zhang et al., 2021a) study how to integrate associated metadata into the label hierarchy for document categorization with weak supervision.
5.3. LLMs as Generators and Annotators
@@
Large language models (LLMs) have demonstrated impressive performance in many downstream tasks and are explored to help low-resource settings by synthesizing data as generators or annotators (Ding et al., 2023). For data generation, few-shot examples (Wang et al., 2021b) or class-conditioned prompts (Meng et al., 2022) are explored for LLM generation and the generated data can be used as pseudo training data to further fine-tune a small model as the final classifier (Ye et al., 2022). Recently, Yu et al. (2023b) proposed an attribute-aware topical text classification method that incorporates ChatGPT to generate topic-dependent attributes and topic-independent attributes to reduce topic ambiguity and increase topic diversity for generation. For data annotation, previous works utilize LLMs for unsupervised annotation (Feng et al., 2021), Chain-of-Thought annotation with explanation generation (He et al., 2024a), and active annotation (Zhang et al., 2023b).
6. Conclusion and Future Work
@@
In this paper, we propose a new method, TELEClass, for the minimally-supervised hierarchical text classification task with two major contributions. First, we enrich the input label taxonomy with LLM-generated and corpus-based class-indicative terms for each class, which can serve as additional features to understand the classes and facilitate classification. Second, we explore the utilization of LLMs in the hierarchical text classification in two directions: data annotation and data creation. On two public datasets, TELEClass can outperform existing baselines substantially, and we further demonstrate its effectiveness through ablation studies. We also conduct a comparative analysis of performance and cost for zero-shot LLM prompting for the hierarchical text classification task.
For future works, first, we plan to generalize TELEClass’s idea of combining LLMs with data-specific knowledge into other low-resource text mining tasks with hierarchical label spaces, such as fine-grained entity typing. Second, in this paper, we mainly focus on acquiring high-quality pseudo labeled data while only utilizing the simplest classifier model and objective. It is worth studying how the proposed method can be further improved with more advanced network structure and noise-robust training objectives. Lastly, we also plan to explore how to extend TELEClass into harder settings like when existing LLMs do not have the knowledge for the initial annotation (e.g., a private domain), a lower-resource scenario where the availability of the unlabeled corpus is limited, or a more complicated label space like an extremely large hierarchical label space with millions of classes.
Acknowledgements.